OpenStudy (anonymous):

find the surface area of in terms of PI A ball has the diamiter of 24

OpenStudy (anonymous):

ok, use the formula that i gave it to you. But remember it's radius in the formula. so you have to convert the diameter to radius. In a diameter, there are 2 radius. Thus, divivde the diameter/2 to get radius. And plug in the information you have into this formula, Surface Area = 4(pi)r^2 Hope you get it... =)

OpenStudy (anonymous):

so it would be 601 ?

OpenStudy (anonymous):

no 616!

OpenStudy (anonymous):

Nope, it's 4(pi)r^2 = 4(pi)(24/2)^2 = 4(144)(pi) = 576 pi

OpenStudy (anonymous):

oh now i see why...i forgot a step!